OpenStudy (anonymous):

help fast 2. 5y + 2 = (10y + 4) (1 point) 5 10 no solution identity 3. 0.5b + 4 = 2(b + 2) (1 point) 0 0.5 no solution identity 4. 8 – (3 + b) = b – 9 (1 point) 7 8 no solution identity

OpenStudy (stormfire1):

#2: 5y + 2 = 10y + 4 2 -4 = 10y-5y 5y = -2 y = -2/5

OpenStudy (stormfire1):

#3 0.5b + 4 = 2(b + 2) 4 = 2b + 4 - .5b 0 = 2b -.5b 1.5b = 0 b = 0

OpenStudy (stormfire1):

#4 8 – (3 + b) = b – 9 8-3-b = b - 9 8-3 = b+ b-9 5 = 2b - 9 2b=-4 b = -2

OpenStudy (anonymous):

Q4 8 – (3 + b) = b – 9 -2b = -14 b = 7

OpenStudy (anonymous):

storm had it until moving the -9

OpenStudy (anonymous):

I LOVE U ALL! thx

OpenStudy (stormfire1):

me with the silly mistake :)